OpenStudy (ella4184):

HELP ASAP PLEASE!!!! will give huge prize!!!! I need 5 fraction to equal exactly 10. but, also one of the fractions has to be either 1/3 and/or 1/4

OpenStudy (anonymous):

i have read this now three times and i have no idea what it means you mean five fractions that add to ten?

OpenStudy (benlindquist):

yes

OpenStudy (anonymous):

ok and one has to be \(\frac{1}{4}\) right?

OpenStudy (benlindquist):

or 1/3

OpenStudy (ella4184):

yes. I am confused too. its part of a project.

OpenStudy (anonymous):

you could do it the easy way

OpenStudy (anonymous):

take \(10-\frac{1}{3}\) then divide the result by 4

OpenStudy (ella4184):

ok, I will try that.

OpenStudy (ella4184):

And then do 119/12 for all five numbers?

OpenStudy (anonymous):

hold on, what is \[10-\frac{1}{3}\]? it is certainly not \(\frac{119}{3}\)

OpenStudy (anonymous):

if you do not know it, say so and i will show you how to get it easily

OpenStudy (ella4184):

oh. i made an error calculating. It is 29/12

OpenStudy (ella4184):

right?

OpenStudy (anonymous):

yes that is correct since \[10-\frac{1}{3}=\frac{29}{3}\] if you add up \[\frac{1}{3}+\frac{29}{12}+\frac{29}{12}+\frac{29}{12}+\frac{29}{12}\] you will get ten

OpenStudy (anonymous):

try it and see

OpenStudy (ella4184):

ok! Thank you! Medal, best answer, and fan!

OpenStudy (anonymous):

thanks !